OpenStudy (zuha):

A circle has an area of approximately 78.5 square units. If the center of the circle is at (2,4). Which of the following points is on the circle?

OpenStudy (zuha):

(-6, 4) (2, -1) (-4, -4) (2, -9)

Nnesha (nnesha):

do you know the standard form equation of the circle ?

OpenStudy (zuha):

Do you mean the distance?

Nnesha (nnesha):

yes same equation for both

OpenStudy (zuha):

Yyeah, I can't understand how to calculate that, I do know the d formula tho ;v;

Nnesha (nnesha):

well the standard form equation of the circle is \[\huge\rm (x-h)^2 +(y-k)=r^2\] where (h,k) is the center and r = radius

Nnesha (nnesha):

use the formula area of a circle \[\huge\rm A= \pi r^2\] to find r (radius )

OpenStudy (zuha):

is (y-k) also supposed to be ^2? So is h=2 and k=4??

Nnesha (nnesha):

yes right forgot sorry about that \[\huge\rm (x-h)^2 +(y-k)^2=r^2\]

Nnesha (nnesha):

yes right h=2 k =4

OpenStudy (zuha):

And I should just try each of the options until one of them is 78.5...?

OpenStudy (zuha):

@Nnesha ?

Nnesha (nnesha):

thanks for tag i was afk alright first you need to find the radius

Nnesha (nnesha):

the given is 78.5 area so substitute a for 78.5 and then solve for r

OpenStudy (zuha):

a?

Nnesha (nnesha):

what do you mean ?

Nnesha (nnesha):

use the formula area of a circle \[\huge\rm A= \pi r^2\] to find radius

OpenStudy (zuha):

Oohhhh 78.5=pir^2 Umm, I assume I divide pi from both sides then?

Nnesha (nnesha):

yes right

OpenStudy (zuha):

25 o.o

OpenStudy (zuha):

So I do 3.14*25^2?

Nnesha (nnesha):

yes r^2 equal 25 yes right forgot sorry about that \[\huge\rm (x-2)^2 +(y-4)^2=25\] now you can substitute (x,y) for each order pair if you get equal sides then that order pair would be the answer

Nnesha (nnesha):

no we need r^2 which is 25 you can check your work 3.14*25^2 iif its equal 78.5 then 25 is correct

OpenStudy (zuha):

It isn't equal to 78.5 .-.

Nnesha (nnesha):

r^2 =25 not r= so dont' take square of 25

Nnesha (nnesha):

\[\huge\rm \frac{ 78.5 }{ 5 }=r^2 ~~~~----> 25=r^2\] to find r you should take square root but we need r^2 for the equation

OpenStudy (zuha):

Ooh Okay, then yeah it's right--so then I just try each of the answers in the slots for x and y until I get the right one?

Nnesha (nnesha):

yep!

OpenStudy (zuha):

Oh okay! I already know the correct answer which is (2,-1) and I tested it and it's right! Thanks so much!! :)

Nnesha (nnesha):

np :=)goood work!